OpenStudy (anonymous):

I have a circuit and I'm trying to find the general solution for a homogeneous equation. My differential equation is: q"+8q'+25q=75 Now, I know that you need to equate it to you so the equation would be: q"+8q'+25q=0 Letting q=e^kt and so on and so forth I've got a gen sol : C1e^(-4t)cos3t+C2e^(-4t)sin3t I've been told that initially the charge on the capacitor and the current in the circuit are zero. So, q(0)=0 and q'(0)=0 How do I find the constants using these initial conditions? I'm looking for guidelines and not the actual answer. Thanks for any help in advance.

OpenStudy (anonymous):

Well, you aren't quite done, after you find the homogenous solution you need to find the particular solution. Which is in a form C. So: \[y_P(x)=C;y'_P(x)=0;y''_P(x)=0\] Plugging this in your see. \[25C=75 \implies C=3\] So your solution would be: \[y(t)=y_H(t)+y_P(t)=c_1e^{-4t}\cos(3t)+c_2e^{-4t}\sin(3t)\] To solve for the constants you simply plug in 0 wherever you see t. Then set your y(t)=q(0)=0. Then differentiate it and do the same thing: y'(t)=q'(0)=0 You'll get a system of equations. Solve for the constants using linear algebra.